OpenStudy (anonymous):

Question 1.1. Find the solution set for the open sentence with the given replacement set. 17 – x = 9, {5, 6, 7, 8} (Points : 1) 5 6 7 8 Question 2.2. Find the solution set for the open sentence with the given replacement set. 2x + 7 = 19, {4, 5, 6, 7} (Points : 1) 4 5 6 7 Question 3.3. Find the solution set for the open sentence with the given replacement set. w2 + 36 = 12w, {0, 2, 4, 6} (Points : 1) 0 2 4 6

OpenStudy (anonymous):

@armygirl99

OpenStudy (cj49):

17-x=9 -x=9-17 x=8

OpenStudy (anonymous):

thank you do u hvae any of the others you could help with

OpenStudy (cj49):

2x=19-17 2x=2 x=1

OpenStudy (cj49):

now can u solve the 3rd one

OpenStudy (anonymous):

x doesnt equal 1 because one isnt an answer

OpenStudy (cj49):

sorryy instead of 7 i thought it was 17 2x+7=19 2x=19-7 2x=12 x=6

OpenStudy (anonymous):

thats better ty for ur help

OpenStudy (anonymous):

could u help me with the rest im not good at this stuff and i have to turn it in

OpenStudy (cj49):

can u substitute the value of w in the 3rd one and try to solve it..

OpenStudy (cj49):

try substituting 6 in the place of w